# Bregman Stochastic Proximal Point Algorithm with Variance Reduction

### v1

@@ 4.1 Convergence Analysis
Suppose that Assumptions 4.5 hold with $\rho>0$. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}\geq\frac{B_{k}}{\rho}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{A_{k}+M_{k}C_{k}}$, $\forall k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$,
$\mathsf{E}[F(\bar{{x}}_{k})-F({\mathsf{x}}_{*})]$ $\leq\frac{(1/\alpha_{0}^{2})\mathsf{E}[D_{h}({\mathsf{x}}_{*},{x}_{0})]+M_{0}\mathsf{E}[\sigma_{0}^{2}]}{\sum_{t=0}^{k-1}(1/\alpha_{t})\left(1-\alpha_{t}(A_{t}+M_{t}C_{t})\right)}+\sum_{t=0}^{k-1}\frac{\mathsf{E}[N_{t}]}{\sum_{t=0}^{k-1}(1/\alpha_{t})\left(1-\alpha_{t}(A_{t}+M_{t}C_{t})\right)},$
@@
Suppose that Assumptions (A.v), 4.5 and (A.iv) are verified with $\rho>0$. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}>\frac{B_{k}}{\rho}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{A_{k}+M_{k}C_{k}}$, $\forall k\in\mathbb{N}$. Set $q_{k}\coloneqq\max\left\{1-\alpha_{k}\gamma_{h}\mu\left(1-\alpha_{k}(A_{k}+M_{k}C_{k})\right),1+\frac{B_{k}}{M_{k}}-\rho\right\}$. Then for all $k\in\mathbb{N}$, $q_{k}\in]0,1[$ and
@@
In both Theorem 4.6 and 4.7, on the right hand sides, the second terms may seem problematic. However, later, for specified variance-reduced algorithms, we will have $N_{k}\leq 0$ for every $k\in\mathbb{N}$. So they can be dropped. Therefore, in terms of order of convergence, these theorems establish the standard sublinear $\mathcal{O}(1/k)$ and linear $\mathcal{O}(q^{k})$ rates for the generic Algorithm 4.1, corresponding, respectively, to convex and relatively strongly convex functions, without requiring a vanishing stepsize. This is an improvement on BSPPA without variance reduction; see Section 5.1. However, the constants in both theorems depend on the sequence $(G_{k})_{k\in\mathbb{N}}$ and this can impede those rates. In the ideal case of Euclidean or quadratic kernel, $G_{k}=1$ [11], and we recover the generic results of Traoré et al. [32].
@@ 5 Instantiation of specific algorithms
In this section, we specialize the generic Algorithm 4.1 with different variance reduction techniques by specifying the term ${e}_{k}$.
@@ 5.1 Bregman Stochastic Proximal Point Algorithm (BSPPA)
Suppose that Assumptions (B.i) and (B.ii) hold with $A_{k}=N_{k}=0$ and $B_{k}\sigma_{k}^{2}\leq\sigma^{2}_{*}\geq 0$ (i.e., $\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}]$ is bounded by $\alpha_{k}^{2}\sigma^{2}_{*}$). Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.1. Then, for $k\geq 1$,
@@
Suppose that Assumptions (A.ii), (B.i), and (B.ii) hold with $\beta>0$ (each $f_{i}$ is relatively strongly convex), $A_{k}=N_{k}=0$ and $B_{k}\sigma_{k}^{2}\leq\sigma^{2}_{*}\geq 0$ (i.e., $\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}]$ is bounded by $\alpha_{k}^{2}\sigma^{2}_{*}$). Let $\alpha_{k}=\alpha>0$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.1. Then, for $k\geq 1$,
@@
In the following, we will present the variance-reduced algorithms. We defer the double-loop, SVRG-style variant of the variance-reduced BSPPA to Appendix A, and focus in the main text on its single-loop counterpart. The first variance-reduced BSPPA that we proposed is using a SAGA-style technique, and we coined it BSAPA. It is the Bregman version of SAPA proposed by Traoré et al. [32].
@@ 5.2 Bregman SAPA (BSAPA)
Suppose that Assumptions (A.iii), 2.1 and (A.v) hold. Let $s\in\mathbb{N}$ and let $({x}_{k})_{k\in[m]}$ be the sequence generated by the inner iteration in Algorithm 5.13. We finally assume that there exists a non-increasing sequence $(G_{k})_{k\in\mathbb{N}}$ such that, for all $i\in[n]$,
@@
$\mathsf{E}_{k}[D_{h}({x}_{k},{z}_{k+1})]$ $\leq 2L\alpha_{k}^{2}G_{k}D_{F}({x}_{k},{\mathsf{x}}_{*})+\alpha_{k}^{2}G_{k}\sigma_{k}^{2}$
$\qquad-\frac{1}{2}D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}]),$ (5.4)
@@
$\sigma_{k}^{2}=2L^{2}$ $\mathsf{E}_{k}\Bigg[D_{h^{*}}\Big(\nabla h({\phi}_{i_{k}}^{k})-\frac{1}{L}(\nabla f_{i_{k}}({\phi}_{i_{k}}^{k})\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({\phi}_{i_{k}}^{k})\Big)\Bigg],$
and
$\mathsf{E}_{k}\left[\sigma_{k+1}^{2}\right]\leq\left(1-\frac{1}{n}\right)\sigma_{k}^{2}+\frac{2L}{n}D_{F}({x}_{k},{\mathsf{x}}_{*}).$ (5.5)
@@
As we stated earlier in Remark 4.4, Lemma 5.8 shows that Assumptions 4.5 are verified with $A_{k}=2LG_{k}$, $B_{k}=G_{k}$, $N_{k}=-(1/2\alpha_{k}^{2})D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])$, $\rho=1/n$ and $C_{k}=C=\frac{2L}{n}$. By just putting this values in Theorems 4.6 and 4.7 we obtain the following two corollaries, respectively.
@@
Let assumptions of Lemma 5.8 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}\geq nG_{k}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.13 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}/n)}$, $\forall k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$,
@@
Let assumptions of Lemma 5.8 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}>nG_{k}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.13 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}/n)}$, $\forall k\in\mathbb{N}$. Set $q_{k}\coloneqq\max\left\{1-\alpha_{k}\gamma_{h}\mu\left(1-2\alpha_{k}L(G_{k}+M_{k}/n)\right),1+\frac{G_{k}}{M_{k}}-\frac{1}{n}\right\}$. Then, for all $k\in\mathbb{N}$, $q_{k}\in]0,1[$,
@@ 5.3 Bregman Loopless SVRP (BLSVRP)
We also proposed BLSVRP, using an L-SVRG-style technique with only one loop. The second loop is replaced by a Bernoulli probability. It generalizes, to the Bregman distance, L-SVRP found for instance in Khaled and Jin [16] and Traoré et al. [32].
@@
Suppose that Assumptions (A.iii), 2.1 and (A.v) hold. Let $s\in\mathbb{N}$ and let $({x}_{k})_{k\in[m]}$ be the sequence generated by the inner iteration in Algorithm 5.13. Assume that there exists a non-increasing sequence $(G_{k})_{k\in\mathbb{N}}$ such that, for all $i\in[n]$,
@@
$\mathsf{E}_{k}[D_{h}({x}_{k},{z}_{k+1})]$ $\leq 2L\alpha_{k}^{2}G_{k}D_{F}({x}_{k},{\mathsf{x}}_{*})+\alpha_{k}^{2}G_{k}\sigma_{k}^{2}$
$\qquad-\frac{1}{2}D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}]),$ (5.9)
@@
$\sigma_{k}^{2}=2L^{2}$ $\mathsf{E}_{k}\Bigg[D_{h^{*}}\Big(\nabla h({u}_{k})-\frac{1}{L}(\nabla f_{i_{k}}({u}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({u}_{k})\Big)\Bigg],$
and
$\mathsf{E}_{k}\big[\sigma_{k+1}^{2}\big]\leq(1-p)\sigma_{k}^{2}+2pLD_{F}({x}_{k},{\mathsf{x}}_{*}).$ (5.10)
@@
Lemma 5.14 shows that Assumptions 4.5 are verified with $A_{k}=2LG_{k}$, $B_{k}=G_{k}$, $N_{k}=-(1/2\alpha_{k}^{2})D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])$, $\rho=p$ and $C_{k}=C=2pL$.
@@
Let assumptions of Lemma 5.14 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}\geq\frac{G_{k}}{p}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.13 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}p)}$, $\forall k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$,
@@
Let assumptions of Lemma 5.14 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}>\frac{G_{k}}{p}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.13 with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}p)}$, $\forall k\in\mathbb{N}$. Set $q_{k}\coloneqq\max\left\{1-\alpha_{k}\gamma_{h}\mu\left(1-2\alpha_{k}L(G_{k}+M_{k}p)\right),1+\frac{G_{k}}{M_{k}}-p\right\}$. Then, for all $k\in\mathbb{N}$, $q_{k}\in]0,1[$,

### v2

@@ 4.1 Convergence Analysis
Suppose that Assumptions 4.5() hold with $\rho>0$. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}\geq\frac{B_{k}}{\rho}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{A_{k}+M_{k}C_{k}}$, $\forall k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$,
@@
Suppose that Assumptions (A.v), 4.5() and (A.iv) are verified with $\rho>0$. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}>\frac{B_{k}}{\rho}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 4.1() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{A_{k}+M_{k}C_{k}}$, $\forall k\in\mathbb{N}$. Set $q_{k}\coloneqq\max\left\{1-\alpha_{k}\gamma_{h}\mu\left(1-\alpha_{k}(A_{k}+M_{k}C_{k})\right),1+\frac{B_{k}}{M_{k}}-\rho\right\}$. Then for all $k\in\mathbb{N}$, $q_{k}\in]0,1[$ and
@@
In both Theorem 4.6() and 4.7(), on the right hand sides, the second terms may seem problematic. However, later, for specified variance-reduced algorithms, we will have $N_{k}\leq 0$ for every $k\in\mathbb{N}$. So they can be dropped. Therefore, in terms of order of convergence, these theorems establish the standard sublinear $\mathcal{O}(1/k)$ and linear $\mathcal{O}(q^{k})$ rates for the generic Algorithm 4.1(), corresponding, respectively, to convex and relatively strongly convex functions, without requiring a vanishing stepsize. This is an improvement on BSPPA without variance reduction; see Section 5.1(). However, the constants in both theorems depend on the sequence $(G_{k})_{k\in\mathbb{N}}$ and this can impede those rates. In the ideal case of Euclidean or quadratic kernel, $G_{k}=1$ [11], and we recover the generic results of Traoré et al. [32].
@@ 5 Instantiation of specific algorithms
In this section, we specialize the generic Algorithm 4.1() with different variance reduction techniques by specifying the term ${e}_{k}$.
@@ 5.1 Bregman Stochastic Proximal Point Algorithm (BSPPA)
Suppose that Assumptions (B.i) and (B.ii) hold with $A_{k}=N_{k}=0$ and $B_{k}\sigma_{k}^{2}\leq\sigma^{2}_{*}\geq 0$ (i.e., $\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}]$ is bounded by $\alpha_{k}^{2}\sigma^{2}_{*}$). Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.1(). Then, for $k\geq 1$,
@@
Suppose that Assumptions (A.ii), (B.i), and (B.ii) hold with $\beta>0$ (each $f_{i}$ is relatively strongly convex), $A_{k}=N_{k}=0$ and $B_{k}\sigma_{k}^{2}\leq\sigma^{2}_{*}\geq 0$ (i.e., $\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}]$ is bounded by $\alpha_{k}^{2}\sigma^{2}_{*}$). Let $\alpha_{k}=\alpha>0$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.1(). Then, for $k\geq 1$,
@@
In the following, we will present the variance-reduced algorithms. We defer the double-loop, SVRG-style variant of the variance-reduced BSPPA to Appendix A(), and focus in the main text on its single-loop counterpart. The first variance-reduced BSPPA that we proposed is using a SAGA-style technique, and we coined it BSAPA. It is the Bregman version of SAPA proposed by Traoré et al. [32].
@@ 5.2 Bregman SAPA (BSAPA)
Suppose that Assumptions (A.iii), 2.1() and (A.v) hold. Let $s\in\mathbb{N}$ and let $({x}_{k})_{k\in[m]}$ be the sequence generated by the inner iteration in Algorithm 5.14(). We finally assume that there exists a non-increasing sequence $(G_{k})_{k\in\mathbb{N}}$ such that, for all $i\in[n]$,
@@
$\mathsf{E}_{k}[D_{h}({x}_{k},{z}_{k+1})]\leq 2L\alpha_{k}^{2}G_{k}D_{F}({x}_{k},{\mathsf{x}}_{*})+\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-\frac{1}{2}D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}]),$ (5.4)
@@
$\sigma_{k}^{2}=2L^{2}$ $\mathsf{E}_{k}\Big[D_{h^{*}}\Big(\nabla h({\phi}_{i_{k}}^{k})-\frac{1}{L}(\nabla f_{i_{k}}({\phi}_{i_{k}}^{k})\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({\phi}_{i_{k}}^{k})\Big)\Big],$
$\text{and }\mathsf{E}_{k}\left[\sigma_{k+1}^{2}\right]\leq\left(1-\frac{1}{n}\right)\sigma_{k}^{2}+\frac{2L}{n}D_{F}({x}_{k},{\mathsf{x}}_{*}).$ (5.5)
@@
As we stated earlier in Remark 4.4(), Lemma 5.8() shows that Assumptions 4.5() are verified with $A_{k}=2LG_{k}$, $B_{k}=G_{k}$, $N_{k}=-(1/2\alpha_{k}^{2})D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])$, $\rho=1/n$ and $C_{k}=C=\frac{2L}{n}$. By just putting this values in Theorems 4.6() and 4.7() we obtain the following two corollaries, respectively.
Assumption 5.10.
For all $k\geq 1$, we assume $G_{k}$ is deterministic and not a random variable. If necessary, we may take $G_{k}=G_{0},\forall k\in\mathbb{N}$. For upper bounds on the gain function $G$, see [11, 14].
@@
Let the assumptions of Lemma 5.8() and Assumption 5.10 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}\geq nG_{k}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.14() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}/n)}$, $\forall k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$,
@@
Let the assumptions of Lemma 5.8() and Assumption 5.10 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}>nG_{k}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.14() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}/n)}$, $\forall k\in\mathbb{N}$. Set $q_{k}\coloneqq\max\left\{1-\alpha_{k}\gamma_{h}\mu\left(1-2\alpha_{k}L(G_{k}+M_{k}/n)\right),1+\frac{G_{k}}{M_{k}}-\frac{1}{n}\right\}$. Then, for all $k\in\mathbb{N}$, $q_{k}\in]0,1[$,
@@ 5.3 Bregman Loopless SVRP (BLSVRP)
We also proposed BLSVRP, using an L-SVRG-style technique with only one loop; see Algorithm 5.14(). The second loop is replaced by a Bernoulli probability. It generalizes, to the Bregman distance, L-SVRP found for instance in Khaled and Jin [16] and Traoré et al. [32].
@@
Suppose that Assumptions (A.iii), 2.1() and (A.v) hold. Let $s\in\mathbb{N}$ and let $({x}_{k})_{k\in[m]}$ be the sequence generated by the inner iteration in Algorithm 5.14(). Assume that there exists a non-increasing sequence $(G_{k})_{k\in\mathbb{N}}$ such that, for all $i\in[n]$,
@@
$\mathsf{E}_{k}[D_{h}({x}_{k},{z}_{k+1})]\leq 2L\alpha_{k}^{2}G_{k}D_{F}({x}_{k},{\mathsf{x}}_{*})+\alpha_{k}^{2}G_{k}\sigma_{k}^{2}-\frac{1}{2}D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}]),$ (5.9)
@@
$\sigma_{k}^{2}=2L^{2}$ $\mathsf{E}_{k}\Big[D_{h^{*}}\Big(\nabla h({u}_{k})-\frac{1}{L}(\nabla f_{i_{k}}({u}_{k})-\nabla f_{i_{k}}({\mathsf{x}}_{*})),\nabla h({u}_{k})\Big)\Big],$
$\text{and }\;\;\mathsf{E}_{k}\big[\sigma_{k+1}^{2}\big]\leq(1-p)\sigma_{k}^{2}+2pLD_{F}({x}_{k},{\mathsf{x}}_{*}).$ (5.10)
@@
Lemma 5.15() shows that Assumptions 4.5() are verified with $A_{k}=2LG_{k}$, $B_{k}=G_{k}$, $N_{k}=-(1/2\alpha_{k}^{2})D_{h^{*}}(\nabla h({x}_{k}),\nabla h({x}_{k})-\mathsf{E}_{k}[\zeta_{k}])$, $\rho=p$ and $C_{k}=C=2pL$.
@@
Let the assumptions of Lemma 5.15() and Assumption 5.10 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}\geq\frac{G_{k}}{p}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.14() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}p)}$, $\forall k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$,
@@
Let the assumptions of Lemma 5.15() and Assumption 5.10 hold. Let $(M_{k})_{k\in\mathbb{N}}$ be a non-increasing positive real-valued sequence such that $M_{k}>\frac{G_{k}}{p}$, $\forall k\in\mathbb{N}$. Suppose also that the sequence $({x}_{k})_{k\in\mathbb{N}}$ is generated by Algorithm 5.14() with $(\alpha_{k})_{k\in\mathbb{N}}$ a non-decreasing positive real-valued sequence such that $\alpha_{k}<\frac{1}{2L(G_{k}+M_{k}p)}$, $\forall k\in\mathbb{N}$. Set $q_{k}\coloneqq\max\left\{1-\alpha_{k}\gamma_{h}\mu\left(1-2\alpha_{k}L(G_{k}+M_{k}p)\right),1+\frac{G_{k}}{M_{k}}-p\right\}$. Then, for all $k\in\mathbb{N}$, $q_{k}\in]0,1[$,